OpenStudy (anonymous):

A paper distributor is shipping poster tubes to an art gallery.The distributor has previosly shipped 560 tubes in 16 boxes.At this rate how many boxes would the distributot need to ship 112 poster tubes ?

OpenStudy (anonymous):

4 boxes

OpenStudy (anonymous):

the last box will not be filled completely.

OpenStudy (anonymous):

ha

OpenStudy (anonymous):

show your work

OpenStudy (anonymous):

560 tubes / 16 boxes = 112 tubes / X boxes solve for X. 560X = 112 * 16 x = 3.2 (so you'd need 4 boxes total)